OpenStudy (anonymous):

What is the equation for the function that is a translation of y=3/x and has asymptotes x=-3 and y=-1? Please show all steps. Will give a medal to best answer!

OpenStudy (anonymous):

@ybarrap @RadEn

OpenStudy (anonymous):

@mathmale?

OpenStudy (mathmale):

Start with the graph of y=1/x. Are you able to graph that? Does the graph have asymptotes? If so, what (where) are they?

OpenStudy (anonymous):

I'm not sure how to graph that. There isn't a graph on the problem. For this assignment I'm supposed to find the error someone else had when they solved the problem, and then I'm supposed to solve it correctly. I don't understand how to do these types of problems.

OpenStudy (anonymous):

I can post the other persons' work for the problem if you would like to see it.

OpenStudy (mathmale):

Yes, post that image if you can.

OpenStudy (anonymous):

And it does say that the asymptotes are x=-3 and y=-1.

OpenStudy (mathmale):

Sarah, I presented you with a simpler example that would be easier to understand. Once you were to grasp the concepts involved in this example, it would be easier for you to udnerstand the actual problem.

OpenStudy (mathmale):

I opened the file you posted, but found nothing but one line of type. :(

OpenStudy (mathmale):

Can't do much until you're able to share an image of the problem you're working on.

OpenStudy (anonymous):

Well, its showing up on my side. So I'll just write out the work that the other person has.

OpenStudy (anonymous):

The problem: What is the equation for the function that is a translation of y=3/x and has asymptotes x=-3 and y=-1? The other persons work that contains an error somewhere: The asymptotes are x=-3 and y=-1, thus h=-3 and k=-1. y=((a)/(x+h))+k y=((3)/(x+(-3)))+(-1) y=((3)/(x-3))-1

OpenStudy (mathmale):

y=((a)/(x+h))+k is incorrect; it should read a y=------ . x-h Substituting h=-3 and k=-1 results in a y = ----- - 1 x+3

OpenStudy (anonymous):

Ooo okay. Then do I just solve from there?

OpenStudy (mathmale):

why don't you actually finish the problem, using the correction I've suggested?